listed below are foot lengths in inches for 11 randomly selected people taken in 1968. find the range, variance, and standard deviation for the given sample data. include apropriate units in the results. are the statistics representative of the current population of all people? 10.1, 9.9, 9.1, 9, 9.9, 10.3, 9.3, 10.4, 9.7, 9.8


1
Expert's answer
2021-10-12T15:09:48-0400

Range = highest - lowest

=10.4 - =1.4

Variance :Following table is obtained for given data :




Mean = "\frac{1}{n} \sum_{i=i}^{n} x_{i}" ="\frac{97.5}{10}=9.75"


Variance:


"\frac{\sum x^2-\frac{(\sum x)^2}{n})}{n-1}\\=\frac{952.7-\frac{(97.5)^2}{10})}{10-1}" = 0.2317


Sd="\sqrt{sd}= \sqrt{0.2317} = 0.4813"